### v2

# On cases where Litt’s game is fair

Abstract
A fair coin is flipped $n$ times, and two finite sequences of heads and tails with the same length are given, say $A$ and $B$. Each time the word $A$ appears in the sequence of fair coin flips, Alice gets a point, and each time the word $B$ appears, Bob gets a point. Who is more likely to win? This puzzle is a slight extension of Litt’s game [8] that recently set Twitter abuzz. We show that Litt’s game is fair for any value of $n$ and any two words that have the same auto-correlation structure by building up a bijection that exchanges Bob and Alice scores; the fact that the inter-correlations do not come into play in this case may come up as a surprise.
1 Introduction and main results
In [8], Litt came up with the following puzzle: a fair coin is tossed $n$ times, Alice gets a point each time the sequence HH appears, while Bob scores a point each time the sequence HT appears (and these sequences may be overlapping). The game may result in a win for Alice, a win for Bob or a tie (the possibilities being exclusive). Who is more likely to win, Alice or Bob ? The perhaps surprising answer (if one judges by the Twitter poll) is that:
  1. 1.
    for any size $n\geqslant 3$, Bob is more likely to win than Alice,
  2. 2.
    for $n$ large, it holds:
    $2\times\big{(}\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})\big{)}\sim\mathbb{P}_{n}(\mbox{Tie})\sim\frac{1}{\sqrt{\pi n}}.$
Several elements of proofs for the first fact quickly emerged on Twitter [13, 14, 17, 18], then more formal proofs of both facts appeared on the arXiv [3, 16]. Nica [10] recorded a YouTube video to introduce the problem to a wider audience. Zeilberger online journal [19] maintains a list of contributions and questions on the problem; among the proofs given, we would like to advertise an early probabilistic proof : in [13], Ramesh builds a “fair majorant” for the score difference of Alice and Bob that consists in a delayed simple random walk; the score difference of Alice and Bob is a deterministic function of this walk, at distance at most 1 below it, see right after the bibliography for more details. A rigorous proof of both facts 1 and 2 is then at an easy reach. Another by-product of his approach is that giving Alice an initial advantage of only one point reverses the statement of fact 1 in a strong sense: for any size $n\geqslant 0$, Alice then has probability strictly greater than $1/2$ of winning. In particular, Alice is more likely to win than Bob. This approach has been detailed in Grimmett [5], who recently rediscovered the arguments of Ramesh.
Considering a contest between HT and TH is also possible, but arguably less interesting: flipping the sequence of tosses, or reading the sequence in the reverse order exchanges Alice and Bob points, which results in a fair game. Our aim in this note is to uncover more hidden symmetries within the words ensuring the game is again tied up.
For $\ell\geqslant 0$, a word of length $\ell$ is a sequence $A=(a_{1},\ldots,a_{\ell})$ in $\{\text{H},\text{T}\}^{\ell}$. For $X_{n}:=(\varepsilon_{k})_{1\leqslant k\leqslant n}$ a finite sequence in $\{\text{H},\text{T}\}^{n}$, we denote by
$N_{A}(X_{n}):=|\{\ell\leqslant k\leqslant n,(\varepsilon_{k-\ell+1},\ldots,\varepsilon_{k})=(a_{1},\ldots,a_{\ell})\}|$
the number of occurrences of the word $A$ in the sequence $X_{n}$. Let $A$ and $B$ be two words of length $\ell$. In the generalized Litt’s game, Alice wins if $N_{A}(X_{n})>N_{B}(X_{n})$, Bob wins if $N_{B}(X_{n})>N_{A}(X_{n})$, and this is a Tie otherwise. A key quantity encoding the intersections of $A$ and $B$ is the correlation of $A$ and $B$, which may be represented as a subset of integers or simply as a number (the base 2 expansion of the subset):
$[A|B]=\sum_{k\in\hbox{{Cor}}(A,B)}2^{k}\quad\text{ where }\quad\mbox{Cor}(A,B)=\{1\leqslant k\leqslant\ell-1,\;(a_{\ell-k+1},\ldots,a_{\ell})=(b_{1},\ldots,b_{k})\}.$
To be more specific, we shall call inter-correlation the correlation of two distinct words (beware the order matters), and auto-correlation the correlation of a word with itself. Our main result in this note is that Litt’s game, which is played with a uniform $X_{n}$, is fair for words $A$ and $B$ with the same auto-correlation, regardless of their inter-correlation.
Theorem 1.
[…]
For instance, the words $A=$ HHTHTH and $B=$ HTTTHH have the same auto-correlation 2 hence Litt’s game played with $A$ and $B$ is fair despite the apparent lack of symmetry between these words.
It is then possible to estimate the quantity in the last equality, or equivalently the probability of a tie $\mathbb{P}_{n}(\text{Tie})$, using a Local Central Limit theorem for sums of weakly dependent random variables. We do not venture into this for the following reason : In a recent breakthrough announced online [11], Nica (together with Janson) pointed to a generic method based on Edgeworth expansions to tackle the case of words $A$ and $B$ with possibly distinct auto-correlations; the sketch of proof, together with some basic moment computations, hints to the following asymptotic estimates: for $A\neq B$, as $n$ gets large, the following asymptotics:
$\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})$ $=\frac{[A|A]-[B|B]}{\sqrt{2^{\ell}+[A|A]+[B|B]-[A|B]-[B|A]}}\cdot\frac{1}{2\sqrt{\pi n}}+o\Big{(}\frac{1}{\sqrt{n}}\Big{)},$
$\mathbb{P}_{n}(\mbox{Tie})$ $=\frac{2^{\ell}}{\sqrt{2^{\ell}+[A|A]+[B|B]-[A|B]-[B|A]}}\cdot\frac{1}{2\sqrt{\pi n}}+o(\frac{1}{\sqrt{n}}\Big{)}$
hold as soon as the denominators on both RHS are non null (the four111(THℓ-1,Hℓ-1T) and its siblings, for which $N_{A}(X_{n})-N_{B}(X_{n})$ belongs to $\{-1,0,1\}$. pair of words giving a null denominator are associated with degenerated cases). Shortly after posting a first version of this work, Segert communicated to us a proof, available online [15] as an addendum to his arXiv paper, of a general statement that entails both estimates. In case $[A|A]=[B|B]$, Theorem 1 refines on the first formula by stating that the quantity is, in fact, identically null for all $n$.
A key feature of the first formula is that $\mathbb{P}_{n}(\mbox{Bob wins})-\mathbb{P}_{n}(\mbox{Alice wins})$ has the sign of $[A|A]-[B|B]$. In light of this observation together with some extensive numerical computations222the conjecture has been checked for all words $A$ and $B$ of size $\leqslant 10$ and for $n\leqslant 35$., we are lead to state the following remarkable conjecture for a fixed length $n$ of the underlying word $X_{n}$. We use $\Delta$ to denote the symmetric difference of two sets.
Conjecture 1.
Assume $[A|A]>[B|B]$. Then, for each $n\geqslant n_{0}:=2\ell-\max\{\hbox{{Cor}}(A)\Delta\hbox{{Cor}}(B)\}$,
[…]
Some comments are in order. First, the restriction to $n\geqslant n_{0}$ is necessary for the first inequality only (the most important one indeed), and we have equality for $n<n_{0}$. Also, the last two inequalities are equivalent (and valid for each $n\geqslant 0$), but we like to phrase the conjecture this way, because of the following vivid interpretation of this set of inequalities: even if Alice looses in the standard version of the game (first inequality), if we were to give Alice one additional point 333beware this is distinct from starting the random word $X_{n}$ with an $A$, then her new winning probability would exceed $1/2$ for each $n$ (last inequality); in particular, Alice would be advantaged over Bob at each subsequent time. Conjecture 1 also suggests that the advantage of Bob versus Alice may be put in bijection with a subset of the Tie event, a key feature of the proof by Ramesh indeed. Conjecture 1 is valid for the original Litt’s game (HH versus HT) by Ramesh’s proof. Also, it is in line with the asymptotics above: to see this, notice the last two inequalities in the conjecture are equivalent to
$|\mathbb{P}_{n}(\text{Alice wins})-\mathbb{P}_{n}(\text{Bob wins})|<\mathbb{P}_{n}(\text{Tie}),$
which may be checked from the conjectured asymptotics because of the inequality $|[A|A]-[B|B]|\leqslant 2^{\ell}-2$ that is true for every pair of words $A$,$B$.
We should also point that the topic of pattern matching and overlaps has been the subject of many investigations, starting with the non-transitive Penney Ante game named after Penney [12] (and famously solved by Conway) in which one is asked to compute the probability that $A$ appears before $B$ in a sequence of fair coin flips, see [6, 7] and the references therein.
Our fixed length conjecture is reminiscent of an important fixed length result in the area of pattern matching : the fact that the number of words $X$ of length $n$ avoiding $A$ is larger than the number of words avoiding $B$ if $[A,A]>[B,B]$ for each $n\geqslant n_{0}$ defined in term of $A$ and $B$ as in Conjecture 1, and with equality before. For $n$ large enough (larger than an indefinite constant), the statement has been stated and proved by Guibas and Odlyzko [6] and independently by Blom [1], while the result for $n\geqslant n_{0}$ has been established by Månnson [9] after a series of works on the topic [4, 2].
2 Proof of Theorem 1
The auto-correlation and inter-correlation of two words are quantities that appear naturally when we look at the probability of one word appearing before another in a sequence of coin flips. The formal definition, which we repeat below, is the following.
Definition 2.
Let $A$, $B$ two words of length $\ell$. We define the indices of inter-correlation of $A$ and $B$ by
$\hbox{{Cor}}(A,B)=\{1\leq k\leq\ell-1,(a_{\ell-k+1},\ldots,a_{\ell})=(b_{1},\ldots,b_{k})\}.$
We write $\hbox{{Cor}}(A)$ to denote $\hbox{{Cor}}(A,A)$. We define the inter-correlation $[A|B]$ of $(A,B)$ by
$[A|B]=\sum_{k\in\hbox{{Cor}}(A,B)}2^{k}$
and the auto-correlation of $A$ as $[A|A]$.
Let us make some remarks about this definition
  • The number $[A|B]$ is not in general equal to $[B|A]$.
  • For all $A,B$ of length $\ell$, $[A|B]\in\llbracket 0;2^{\ell}-2\rrbracket$.
  • For every words $A,B,C,D$ of length $\ell$, $[A|B]=[C|D]$ if and only if $\hbox{{Cor}}(A,B)=\hbox{{Cor}}(C,D)$.
[…]
We introduce some additional notation. If $C,D$ are two words, $CD$ will be the concatenation of $C$ and $D$. Besides, for two words $C,D$ of length $\ell$ and $m\in\hbox{{Cor}}(C,D)$, we denote $C^{m}D$ the word of length $2\ell-m$ beginning by $C$ and ending by $D$. For example, if $C=\underline{\text{HTTHTH}}$, $D=\overline{\text{THTHHH}}$, we have $C^{2}D=\underline{\text{HTTH}}\underline{\overline{\mathbf{\text{TH}}}}\overline{\text{THHH}}$. We extend this notation to $k$ words $C_{1},\ldots,C_{k}$ of length $\ell$ and $m_{1},\ldots,m_{k}\geqslant 1$ in the obvious way: if $m_{i}\in\hbox{{Cor}}(C_{i},C_{i+1})$ for each $i$, we define the word of length $k\ell-\sum m_{i}$
$Y=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}C_{3}\mbox{}^{m_{3}}\ldots C_{k-1}\mbox{}^{m_{k-1}}C_{k}.$ (1)
Definition 3.
Let $A,B$ be two words of length $\ell$. We call overlap of $A$ and $B$ a word $Y$ of the form (1) where $C_{1},\ldots,C_{k}\in\{A,B\}^{k}$. We denote by $\mathcal{E}(A,B)$ the set of all overlaps of $A$ and $B$.
Note that do not accept $m_{i}=0$ in (1). In particular, the concatenation $Y=AB$ of $A$ and $B$ may not be in $\mathcal{E}(A,B)$. We notice also that, for $Y\in\mathcal{E}(A,B)$, the expression of $Y$ in the form given by (1) may be not unique. We call the maximal decomposition of $Y$ the one such that, if $Y=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}C_{3}\mbox{}^{m_{3}}\ldots C_{k-1}\mbox{}^{m_{k-1}}C_{k}$, we have
$N_{A}(Y)=|\{1\leq i\leq k,C_{i}=A\}|\qquad N_{B}(Y)=|\{1\leq i\leq k,C_{i}=B\}|.$
We have the following decomposition of a word.
Definition 4.
Let $Y$ a word. There exists a unique way to write $Y$ in the form
$Y=X_{0}E_{1}X_{1}E_{2}\ldots X_{k-1}E_{k}X_{k}$ (2)
such that
  • the words $E_{1},\ldots,E_{k}$ belong to $\mathcal{E}(A,B)$.
  • For all $i\in\llbracket 0,k\rrbracket$, neither $A$ nor $B$ appears in $X_{i}$ (word $X_{i}$ may be empty).
  • $N_{A}(Y)=\sum_{i=1}^{k}N_{A}(E_{i})$ and $N_{B}(Y)=\sum_{i=1}^{k}N_{B}(E_{i})$.
We call pattern of the word $Y$ with respect to $A$ and $B$ the words $(E_{1},\ldots,E_{k})$ which appears in (2) and write $\textbf{Patt}_{A,B}(Y)=(E_{1},E_{2},\ldots,E_{k})$.
Proposition 1.
Let $A,B$ be two words with the same auto-correlation, and let $\phi:\mathcal{E}(A,B)\to\mathcal{E}(A,B)$ be defined in the following way. If $Y:=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}\ldots C_{k-1}\mbox{}^{m_{k-1}}C_{k}$ with $C_{i}\in\{A,B\}$, we set
$\phi(Y)=\bar{C}_{k}\mbox{}^{m_{k-1}}\bar{C}_{k-1}\mbox{}^{m_{k-2}}\ldots\bar{C}_{2}\mbox{}^{m_{1}}\bar{C}_{1}$
where $\bar{C}_{i}=A$ if $C_{i}=B$ and $\bar{C}_{i}=B$ if $C_{i}=A$. Then $\phi$ is well-defined, it is independent of the decomposition chosen for $Y$, it is an involution and $\phi(Y)$ has the same length as $Y$. Moreover, we have $(N_{A}(Y),N_{B}(Y))=(N_{B}(\phi(Y)),N_{A}(\phi(Y)))$.
Proof.
We first prove that $\phi(Y)$ is well defined i.e. if $m_{i}\in\text{Cor}(C_{i},C_{i+1})$, then $m_{i}\in\text{Cor}(\bar{C}_{i+1},\bar{C}_{i})$. Recall that we assume that $A,B$ have the same auto-correlation, i.e. $\text{Cor}(A)=\text{Cor}(B)$. We have two cases:
  • Either $C_{i}=C_{i+1}$, for example $C_{i}=A$. Then $\bar{C}_{i}=\bar{C}_{i+1}=B$. Hence, we have $\text{Cor}(C_{i},C_{i+1})=\text{Cor}(A)=\text{Cor}(B)=\text{Cor}(\bar{C}_{i+1},\bar{C}_{i})$.
  • Or $C_{i}\neq C_{i+1}$, for example $(C_{i},C_{i+1})=(A,B)$. Then $(\bar{C}_{i+1},\bar{C}_{i})$ is also equal to $(A,B)$ and so $\text{Cor}(C_{i},C_{i+1})=\text{Cor}(\bar{C}_{i+1},\bar{C}_{i})$.
Note that $\phi$ does not depend of the decomposition chosen for $Y$. To justify this claim, it is enough to consider the case of words $Y$ with two decompositions $Y=C_{1}\mbox{}^{m_{1}}C_{2}\mbox{}^{m_{2}}C_{3}$ and $Y=C_{1}\mbox{}^{m}C_{3}$. Note that $|Y|=2\ell-m=3\ell-m_{1}-m_{2}$. Applying $\phi$ with the first decomposition we get $\phi(Y)=\bar{C}_{3}\mbox{}^{m_{2}}\bar{C}_{2}\mbox{}^{m_{1}}\bar{C}_{1}$ and since $|\phi(Y)|=|Y|<2\ell$, necessarily, $\bar{C}_{3}$ and $\bar{C}_{1}$ overlap in the writing of $\phi(Y)$ and thus we also have $\phi(Y)=\bar{C}_{3}\mbox{}^{m}\bar{C}_{1}$. The fact that $\phi$ is an involution is clear.
We write now $Y$ with its maximal decomposition. By construction, we directly get that $N_{A}(\phi(Y))\geqslant N_{B}(Y)$ and $N_{B}(\phi(Y))\geqslant N_{A}(Y)$. We claim that, by maximality of the decomposition, there is equality in these two inequalities. Indeed, consider the case of words $Y$ whose maximal decomposition consist in two words: $C_{1}\mbox{}^{m}C_{2}$. Among those words $Y$, only the words where $C_{1}=C_{2}$ have to be considered. If $Y_{0}=C_{1}\mbox{}^{m}C_{1}$, it holds $\phi(Y_{0})=\bar{C}_{1}\mbox{}^{m}\bar{C}_{1}$. Now, assume that $\phi(Y_{0})=\bar{C}_{1}\mbox{}^{m_{1}}C_{3}\mbox{}^{m_{2}}\bar{C}_{1}$ for some $C_{3}\in\{A,B\}$. The map $\phi$ being independent of the decomposition, we get $Y_{0}=\phi(\phi(Y_{0}))=C_{1}\mbox{}^{m_{1}}\bar{C_{3}}\mbox{}^{m_{2}}C_{1}$, which has a distinct count of the $\bar{C_{3}}$-word; this negates the fact that $Y_{0}=C_{1}\mbox{}^{m}C_{1}$ is the maximal decomposition in the first place. ∎
The function $\phi$ defined in the previous proposition can be extended to a function on patterns. For $(E_{1},\ldots,E_{k})\in\mathcal{E}(A,B)^{k}$, we set
$\phi(E_{1},\ldots,E_{k})=(\phi(E_{k}),\ldots,\phi(E_{1}))$
which defines an involution on $\mathcal{E}(A,B)^{k}$. Given a pattern $M=(E_{1},\ldots,E_{k})\in\mathcal{E}(A,B)^{k}$ and $n\geqslant 1$, we define $L_{M}(n)$ as the number of words of length $n$ with pattern $M$:
$L_{M}(n)=|\{\hbox{word $Y$},\;|Y|=n\hbox{ and }\textbf{Patt}_{A,B}(Y)=M\}|.$
To prove Theorem 1, it is sufficient to show that, for any pattern $M$ and any $n\geqslant 1$, we have
$L_{M}(n)=L_{\phi(M)}(n).$ (3)
Indeed, if $Y$ is a word such that $\textbf{Patt}_{A,B}(Y)=M$ and $Z$ is a word such that $\textbf{Patt}_{A,B}(Z)=\phi(M)$, we have
$(N_{A}(Y),N_{B}(Y))=(\sum_{i=1}^{k}\!N_{A}(E_{i}),\sum_{i=1}^{k}\!N_{B}(E_{i}))=(\sum_{i=1}^{k}\!N_{B}(\phi(E_{i})),\sum_{i=1}^{k}\!N_{A}(\phi(E_{i})))=(N_{B}(Z),N_{A}(Z)).$
In view of Equality (3), we conclude that
$P\big{(}(N_{A}(X_{n}),N_{B}(X_{n}))=(a,b)\big{)}\;=\;\frac{1}{2^{n}}\sum_{\begin{subarray}{c}\hbox{\scriptsize{pattern $M$ s.t.}}\\ \hbox{\tiny{$(N_{A}(M),N_{B}(M))=(a,b)$}}\end{subarray}}L_{M}(n)\;$ $=$ $\;\frac{1}{2^{n}}\sum_{\begin{subarray}{c}\hbox{\scriptsize{pattern $M$ s.t.}}\\ \hbox{\tiny{$(N_{A}(M),N_{B}(M))=(a,b)$}}\end{subarray}}L_{\phi(M)}(n)$
$=$ $P\big{(}(N_{B}(X_{n}),N_{A}(X_{n}))=(a,b)\big{)}.$
In order to establish (3), we prove the more slightly more precise result:
Lemma 1.
For any pattern $M=(E_{1},\ldots,E_{k})$, for any $I=(i_{0},\ldots,i_{k})$, set
$L^{M}(I)=|\{Y=X_{0}E_{1}X_{1}\ldots E_{k}X_{k}:\textbf{Patt}_{A,B}(Y)=M;\forall j,\,|X_{j}|=i_{j}\}|.$
Then
$|L^{M}(I)|=|L^{\phi(M)}(I^{\prime})|$
where $I^{\prime}=(i_{k},\ldots,i_{0})$.
Proof.
Let us remark that we have
$L^{M}(I)=L^{E_{1}}(i_{0},0)\left(\prod_{j=1}^{k-1}L^{(E_{j},E_{j+1})}(0,i_{j},0)\right)L^{E_{k}}(0,i_{k})$
and
$L^{\phi(M)}(I^{\prime})=L^{\phi(E_{k})}(i_{k},0)\left(\prod_{j=1}^{k-1}L^{(\phi(E_{j+1}),\phi(E_{j}))}(0,i_{j},0)\right)L^{\phi(E_{1})}(0,i_{0}).$
Moreover, the value of $L^{(E_{j},E_{j+1})}(0,i_{j},0)$ only depends on $i_{j}$, on the word ending $E_{j}$ and on the word beginning $E_{j+1}$. For example, if $E_{j}$ ends with an $A$ and $E_{j+1}$ starts with a $B$, we have
$L^{(E_{j},E_{j+1})}(0,i_{j},0)=L^{(A,B)}(0,i_{j},0).$
Now, if $E_{j}$ ends with an $A$ and $E_{j+1}$ starts with a $B$, then $\phi(E_{j+1})$ ends with an $A$ and $\phi(E_{j})$ starts with an $B$. Thus, in this case, we directly get
$L^{(\phi(E_{j+1}),\phi(E_{j}))}(0,i_{j},0)=L^{(E_{j},E_{j+1})}(0,i_{j},0)=L^{(A,B)}(0,i_{j},0).$
The situation is more involved when $E_{j}$ ends with the same word than $E_{j+1}$ starts with, let say the word $A$. Then indeed
$L^{(\phi(E_{j+1}),\phi(E_{j}))}(0,i_{j},0)=L^{(B,B)}(0,i_{j},0)\mbox{ while }L^{(E_{j},E_{j+1})}(0,i_{j},0)=L^{(A,A)}(0,i_{j},0).$
Besides, if $E_{1}$ starts with $A$, then $\phi(E_{1})$ ends with a $B$, and we have
$L^{\phi(E_{1})}(0,i_{0})=L^{(B)}(0,i_{0})\mbox{ while }L^{E_{1}}(i_{0},0)=L^{(A)}(i_{0},0),$
and a similar assertion holds for $L^{E_{k}}(0,i_{k})$ and $L^{\phi(E_{k})}(i_{k},0)$. Combining all these remarks, we see that Lemma 1 will be proved as soon as we establish the following equality: for each $i\geqslant 0$,
$L^{(A,A)}(0,i,0)=L^{(B,B)}(0,i,0)\qquad\mbox{and}\qquad L^{(B)}(0,i)=L^{(A)}(i,0).$ (4)
We prove (4) by induction on $i$. The proposition clearly holds for $i=0$. Assume it holds for $k\leq i-1$. Thus, for any pattern $M$, if $I=(i_{0},\ldots,i_{k})$ with $i_{j}<i$ for all $j$, we get
$|L^{M}(I)|=|L^{\phi(M)}(I^{\prime})|.$
Let us now prove that $L^{(B)}(0,i)=L^{(A)}(i,0)$. Writing $|I|=\sum_{j=0}^{k}i_{j}$ if $I=(i_{0},\ldots,i_{k})$ and $|M|=\sum_{i=1}^{k}|E_{i}|$ for the length of the pattern $M=(E_{1},\ldots,E_{k})$, we find that
$L^{(A)}(i,0)=$ $=$ |$|\{\hbox{words $XA$}:\;|X|=i\hbox{ and }\textbf{Patt}_{A,B}(XA)=A\}|$
$=$ $2^{i}-\sum_{k\geqslant 1}\hskip 2.84544pt\sum_{\begin{subarray}{c}M=(E_{1},\ldots,E_{k})\neq(A)\\ \hbox{\scriptsize{$E_{k}\mbox{ ends with }A$}}\end{subarray}}\hskip 8.5359pt\sum_{\begin{subarray}{c}I=(i_{0},\ldots,i_{k-1},0)\\ |I|+|M|=i+|A|\end{subarray}}|L^{M}(I)|$
$=$ $2^{i}-\sum_{k\geqslant 1}\hskip 2.84544pt\sum_{\begin{subarray}{c}M=(E_{1},\ldots,E_{k})\neq(A)\\ \hbox{\scriptsize{$E_{k}\mbox{ ends with }A$}}\end{subarray}}\hskip 8.5359pt\sum_{\begin{subarray}{c}I=(i_{0},\ldots,i_{k-1},0)\\ |I|+|M|=i+|A|\end{subarray}}|L^{\phi(M)}(I^{\prime})|.$
At this point, we use the recurrence assumption noticing that $|I|<i$ since $|M|>|A|$. Recalling that if $M$ ends with $A$, then $\phi(M)$ starts with $B$, we see that the last line is also equal to
$2^{i}-\sum_{k\geqslant 1}\hskip 2.84544pt\sum_{\begin{subarray}{c}M=(E_{1},\ldots,E_{k})\neq(B)\\ \hbox{\scriptsize{$E_{1}\mbox{ starts with }B$}}\end{subarray}}\hskip 8.5359pt\sum_{\begin{subarray}{c}I=(0,i_{1},\ldots,i_{k})\\ |I|+|M|=i+|B|\end{subarray}}|L^{M}(I)|\;=\;L^{(B)}(0,i).$
The equality $L^{(A,A)}(0,i,0)=L^{(B,B)}(0,i,0)$ is proved with a similar argument. ∎